Freedom Bank has received a loan application from a private company Automatic Corporation. An abbreviated portion of the financial information provided by the company is shown below:

       


Total assets                                                          $1,260,000

EBIT                                                                         158,000

Net working capital                                                    84,000

Book value of equity                                                 360,000

Accumulated retained earnings                                320,000

Total liabilities                                                       1,140,000

 

 

What is the Z-score for Automatic Corporation? What does the score predict?       



1
Expert's answer
2022-03-02T10:26:03-0500

"Z=\frac{84000}{1 260 000}+\frac{320 000}{1 260 000}+\frac{158 000}{1260 000}+\frac{360 000}{1140 000}=0.067+0.254+0.125+0.316=0.762"


1.1 and less - "Red" zone, there is a possibility of bankruptcy of the enterprise.